# Methods and Trends in Detecting Generated Images: A Comprehensive Review

### v1

@@
Abstract.
The proliferation of generative models, such as Generative Adversarial Networks (GANs), Diffusion Models, and Variational Autoencoders (VAEs), has enabled the synthesis of high-quality multimedia data. However, these advancements have also raised significant concerns regarding adversarial attacks, unethical usage, and societal harm. Recognizing these challenges, researchers have increasingly focused on developing methodologies to detect synthesized data effectively, aiming to mitigate potential risks. Prior reviews have primarily focused on deepfake detection and often lack coverage of recent advancements in synthetic image detection, particularly methods leveraging multimodal frameworks for improved forensic analysis. To address this gap, the present survey provides a comprehensive review of state-of-the-art methods for detecting and classifying synthetic images generated by advanced generative AI models. This review systematically examines core detection methodologies, identifies commonalities among approaches, and categorizes them into meaningful taxonomies. Furthermore, given the crucial role of large-scale datasets in this field, we present an overview of publicly available datasets that facilitate further research and benchmarking in synthetic data detection.
Forensic Image Detection, Fingerprint Analysis, Generative Adversarial Networks (GANs), Diffusion Models, Frequency Domain Analysis, Vision-Language Models (VLMs), Patch-Based Detection, Cross-Model Generalization, Cross-Domain Detection, Cross-Scene Analysis, Synthetic Image Identification
1. Introduction
The advent of advanced generative models has enabled the creation of highly realistic synthetic images. These images are synthesized through various approaches, including conditional methods such as image-to-image translation and text-to-image translation, as well as unconditional generation. The present survey focuses on the detection of fully synthesized images, distinguishing them from subtle manipulations or partial modifications. While prior survey have addressed the detection of image manipulations achieved through traditional computer-based techniques (Sencar and Memon, 2013), image editing tools (Zheng et al., 2019), and deepfake methodologies (Wang et al., 2024), our focus diverges from these areas. Specifically, we examine detection methods designed to identify images synthesized by state-of-the-art generative models (Nichol et al., 2021; Ramesh et al., 2021; Rombach et al., 2022; Wukong, 2022), including recent approaches that incorporate multimodal techniques (Radford et al., 2021) for enhanced detection accuracy. Given the increasing relevance of such methods in forensic image analysis, it is crucial to explore their long-term applicability and effectiveness in generative image detection.
The detection of images generated by both conditional and unconditional generative methods has emerged as a rapidly growing research area. However, despite this increasing interest, a comprehensive review consolidating and systematically categorizing existing detection methodologies remains lacking. This survey addresses this gap by providing a structured analysis of detection approaches, classifying them based on their underlying techniques and contributions.
Figure 1. Illustration of Simplified Architecture of GAN, Diffusion, VAE and Autoregressive Generative Models.
@@ 1. Introduction
Diagram illustrating the architectures of various generative models
Generative Adversarial Networks (GANs)
GANs were first proposed by Goodfellow et al. (Goodfellow et al., 2014). The architecture consists of two neural networks, a generator $G$ and a discriminator $D$, that compete in an adversarial game, optimizing the adversarial loss:
(1) $\min_{G}\max_{D}\mathbb{E}_{x\sim p_{data}(x)}[\log D(x)]+\mathbb{E}_{z\sim p_{z}(z)}[\log(1-D(G(z)))].$
This innovative approach introduced adversarial training, boosting the quality of generated data. Despite their success, GANs face challenges such as mode collapse and instability during training.
Diffusion Models
Diffusion models were first introduced by Sohl-Dickstein et al. (Sohl-Dickstein et al., 2015) and later demonstrated in practice by Ho et al. (Ho et al., 2020). These models progressively convert an initial Gaussian noise distribution into a structured data distribution via a bidirectional diffusion process. The reverse process is guided by the following equation:
(2) $\mathbf{x}_{t-1}=\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\alpha_{t}}}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})+\sqrt{\beta_{t}}\mathbf{z}_{t},$
where $\beta_{t}$ and $\alpha_{t}$ are variance scaling parameters. Diffusion models have achieved state-of-the-art results in image generation and have played a pivotal role in advancing text-to-image generation. However, they remain computationally intensive due to the iterative denoising process.
Variational Autoencoders (VAEs)
VAEs were proposed by Kingma and Welling (Kingma, 2013). VAEs learn a probabilistic mapping between data and a latent space by maximizing the Evidence Lower Bound (ELBO):
(3) $\mathcal{L}=\mathbb{E}_{q_{\phi}(z|x)}[\log p_{\theta}(x|z)]-D_{KL}(q_{\phi}(z|x)||p(z)).$
This framework enables efficient data sampling (encoding) and reconstruction (decoding). Despite their innovative probabilistic approach, VAEs often generate blurry images compared to GANs due to their reliance on pixel-wise reconstruction loss.
Autoregressive Models
Autoregressive models for image generation were first introduced in PixelRNN (Van Den Oord et al., 2016) and PixelCNN (Van den Oord et al., 2016) by Oord et al. These models factorize the joint probability distribution of an image as a product of conditional probabilities, generating pixels sequentially:
(4) $p(\mathbf{x})=\prod_{i=1}^{N}p(x_{i}|x_{<i}),$
where each pixel $x_{i}$ is predicted based on previously generated pixels. PixelRNN uses recurrent layers to model spatial dependencies, while PixelCNN employs masked convolutions to improve parallelization. Recently, Transformer-based autoregressive models(Parmar et al., 2018) have further improved scalability by leveraging self-attention mechanisms. Autoregressive models achieve high-quality image synthesis with explicit likelihood estimation but suffer from slow generation due to their sequential nature.
@@
While these four generative model families, GANs, Diffusion Models, VAEs, and Autoregressive, have dominated the landscape of image generation, other methods such as Normalizing Flows (Rezende and Mohamed, 2015), have also made significant contributions and deserve attention. With these advancements, several state-of-the-art generative models have been made publicly available alongside commercially accessible tools such as Adobe Firefly (Adobe Inc., 2024), MidJourney (MidJourney, 2024), DALL·E 3 (OpenAI, 2024), and Imagen 3 (Google DeepMind, 2024).
@@
Although generative models have enabled significant advancements in image synthesis, their accessibility introduces critical concerns related to misinformation, privacy, and security. The ability to generate highly realistic AI-synthesized images raises ethical and security risks, including deceptive media (Sandrini and Somogyi, 2023), identity fraud (Zhang et al., 2025), and geopolitical manipulation (Mylrea, 2025). Consequently, forensic detection has become an essential area of research, aiming to reliably distinguish AI-generated imagery from real-world photographs. Despite significant progress, existing detection techniques face challenges in generalizability, robustness, and scalability across evolving generative architectures. This survey provides a comprehensive review of image detection methodologies, summarizing key advancements, identifying existing limitations, and outlining future directions. By examining the evolving landscape of adversarial threats in generative AI, this study highlights critical challenges and opportunities, reinforcing the need for continued research in forensic detection.
2. Reviews
@@
In the present survey of detection methods for AI-generated images, we categorize them into six distinct groups, with the final category comprising commercial methods. For each group, we discuss the core proposal of each method in ascending order based on publication date. At the end of each subsection, we present a table summarizing whether the experimental evaluations of the methods satisfy three key criteria, described as follows:
@@
3. Cross-Scene: This criterion assesses whether a method’s performance was tested across datasets with distinct data distributions. For instance, a detector trained on images from the ImageNet (Russakovsky et al., 2015) dataset and evaluated on images from the LSUN-Bedroom (Yu et al., 2015) dataset satisfies the cross-scene requirement. Importantly, methods that meet the cross-scene criterion typically also satisfy the cross-category requirement, although the converse is not necessarily true.
@@
Figure 2. Categorization of Detection Methods Based on Core Architecture and Methodological Proposals. The figure illustrates the division of detection methods into categories. Each category is highlighted using blue-colored rectangles. Some methods are connected to multiple categories, shown using light-red highlights and arrows pointing to the respective sub-categories.
[…]
Diagram illustrating the categories.
2.1. Spatial Domain Analysis / Spatial Feature Analysis Methods
@@
2.1.1. Co-occurrence Matrices On Detecting GAN Generated Images by Nataraj et al.: 2019
Nataraj et al. (Nataraj et al., 2019) proposed a GAN-generated image detection method using co-occurrence matrices extracted from RGB channels. Unlike forensic techniques that used such matrices, their approach bypasses residual computations and directly processes pixel co-occurrence matrices using a deep convolutional neural network (CNN). The method computes co-occurrence matrices on the red, green, and blue channels of an image, capturing statistical anomalies introduced by GAN-generated images. These matrices are structured into tensors, which are subsequently processed through a custom CNN. The network consists of convolutional layers with ReLU activations, max pooling operations, and fully connected layers, concluding with a sigmoid activation for binary classification. The evaluation was conducted on two diverse datasets: CycleGAN (Zhu et al., 2017) (unpaired image-to-image translations) and StarGAN (Choi et al., 2018) (facial attribute manipulations).
2.1.2. CNN-Generated Images Are Surprisingly Easy to Spot: Wang et al. (2020)
Wang et al. (Wang et al., 2020b) conducted a pivotal study demonstrating that a detection method trained on images generated by a single generative model, particularly a GAN, can generalize to detect synthetic images from a variety of unseen CNN-based generative models. This finding challenges the previously established view that cross-model generalization is inherently difficult for forensic classifiers. To evaluate this, the authors trained ProGAN (Karras, 2017) on the LSUN dataset (Yu et al., 2015), producing a dataset of 720K images for training and 4K for validation, with an equal split between real and generated images. They employed a ResNet-50 (He et al., 2016), pretrained on ImageNet, as a binary classifier. Robust feature learning was facilitated through various data augmentation strategies, including: (a) no augmentation, (b) Gaussian blur, (c) JPEG compression, and (d) combinations of both with varying probabilities (50% and 10%). The trained classifier demonstrated strong generalization capabilities, successfully detecting images synthesized by other prominent generative models such as StyleGAN (Karras et al., 2019), BigGAN (Brock, 2018a), CycleGAN (Zhu et al., 2017), StarGAN (Choi et al., 2018), and GauGAN (Park et al., 2019). This work highlights the existence of common artifacts across CNN-generated images, suggesting that classifiers can leverage these shared patterns for generalizable detection across different architectures and tasks.
2.1.3. Detection and Attribution of GAN Images: Goebel et al. (2020)
Goebel et al. (Goebel et al., 2020) introduced a comprehensive framework to detect, attribute, and localize GAN-generated images through co-occurrence matrix analysis and deep learning. This method builds on insights from steganalysis, leveraging pixel-level co-occurrence features to identify artifacts introduced during image synthesis. The approach begins by computing co-occurrence matrices from the RGB channels of the input image in four orientations: horizontal, vertical, diagonal, and anti-diagonal. Each co-occurrence matrix captures a 256×256 histogram of pixel value pairs, normalized and stacked into a tensor of size $256\times 256\times 12$. The matrices are defined as:
@@
These features are then processed using a modified XceptionNet (Chollet, 2017), designed for three key tasks: 1. Binary Detection: Classifying images as real or GAN-generated. 2. Multi-Class Attribution: Identifying the GAN architecture (e.g., ProGAN, CycleGAN). 3. Localization: Generating heatmaps to identify manipulated regions through patch-based analysis.
Extensive experiments conducted on over 2.76 million images demonstrated the effectiveness of this method across various GAN models, including ProGAN, StyleGAN, CycleGAN, StarGAN, and SPADE/GauGAN. Additionally, t-SNE (Van der Maaten and Hinton, 2008) visualizations showed a clear separation between real and GAN-generated images, reinforcing the model’s interpretability and robustness against varying JPEG compression levels and patch sizes. This framework advances GAN forensics by integrating detection, attribution, and localization into a unified pipeline.
2.1.4. Estimating Artifact Similarity with Representation Learning: Li et al. (2022), GASE-Net
Li et al. (Li et al., 2022) introduced GASE-Net, a framework designed to detect GAN-generated images by estimating artifact similarity. This method addresses challenges in cross-domain generalization and robustness against post-processing, using a two-stage approach: representation learning and representation comparison. In the representation learning stage, a ResNet-50 backbone is modified with instance normalization (IN) applied to the shallow layers to enhance feature extraction by filtering out instance-specific biases. This ensures that the learned representations remain invariant across different domains while retaining category-level distinctions. Feature maps from reference images are aggregated element-wise to form domain prototypes, which serve as robust representations of GAN or pristine image domains.
@@
2.1.5. Local Intrinsic Dimensionality Analysis: Lorenz et al. (2023), AdaptedMultiLID
Lorenz et al. (Lorenz et al., 2023) proposed a framework utilizing the Multi-Local Intrinsic Dimensionality (multiLID) method for detecting diffusion-generated images. This approach builds upon the earlier work (Lorenz et al., 2022) and demonstrates strong performance in distinguishing both diffusion and GAN-generated images. The method starts by extracting low-dimensional feature maps using an untrained ResNet18 model. MultiLID is then computed to capture local growth rates of feature densities within the latent space. The multiLID feature vector for each sample $x$ is defined as:
@@
2.1.6. AI-Generated Image Detection: Baraheem et al. (2023)
Baraheem et al. (Baraheem and Nguyen, 2023) proposed a framework to detect GAN-generated images using transfer learning on pretrained classifiers. The authors compiled a diverse dataset called Real and Synthetic Images (RSI), consisting of 48,000 images synthesized by 12 GAN architectures across tasks such as image-to-image, sketch-to-image, and text-to-image generation. EfficientNetB4 (Tan and Le, 2019) achieved the best detection performance after fine-tuning on the RSI dataset. The model’s architecture was modified by replacing the classifier head with layers for global average pooling, batch normalization, ReLU activation, dropout, and a sigmoid output. The training utilized the Adam optimizer with a batch size of 64, an initial learning rate of 0.001, and data augmentation techniques such as horizontal flips. To facilitate model explainability, the authors incorporated four Class Activation Map (CAM) methods, GradCAM (Selvaraju et al., 2017), AblationCAM (Ramaswamy et al., 2020), LayerCAM (Jiang et al., 2021), and Faster ScoreCAM (Wang et al., 2020a), to visualize the discriminative regions influencing classification decisions.
2.1.7. Diffusion Reconstruction Error (DIRE): Wang et al. (2023)
Wang et al. (Wang et al., 2023a) proposed DIffusion Reconstruction Error (DIRE), a novel method to detect diffusion-generated images by leveraging reconstruction errors from pre-trained diffusion models. The method addresses the limitations of previous detectors, which struggled to generalize across different diffusion models. The key idea is that diffusion-generated images can be reconstructed more accurately by the pre-trained DDIM model (Song et al., 2020) than real images. DIRE is defined as the $L_{1}$-norm of the difference between the input image $x_{0}$ and its reconstructed counterpart $x^{\prime}_{0}$:
@@
The process involves applying forward noise to the input image and then performing a reverse denoising process to generate the reconstruction. A ResNet-50 classifier is trained using binary cross-entropy loss on these DIRE representations to distinguish between real and generated images. The authors demonstrated that DIRE achieves state-of-the-art performance on their proposed DiffusionForensics dataset, which includes images generated by various diffusion models across multiple domains (e.g., LSUN-Bedroom (Yu et al., 2015), ImageNet (Russakovsky et al., 2015), and CelebA-HQ (Karras, 2017)). Extensive experiments showed that DIRE not only excels in detecting diffusion-generated images but also generalizes well to unseen models and maintains robustness under perturbations like Gaussian blur and JPEG compression.
2.1.8. Classification and Explainable Identification: Bird and Lotfi (2024)
Bird and Lotfi (Bird and Lotfi, 2024) introduced a framework for detecting AI-generated images using a large-scale dataset, CIFAKE, which is detailed in the datasets section. Their classification approach employs a Convolutional Neural Network (CNN) that processes images through stacked convolutional and pooling layers, followed by fully connected layers with a final sigmoid activation for binary classification. As seen in (Baraheem and Nguyen, 2023), the study emphasizes explainability by implementing Gradient Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2017), which generates heatmaps to highlight regions influencing the model’s decisions. Grad-CAM computes importance weights $\alpha_{k}$ for each feature map $A_{k}$, producing a visual explanation as:

### v2

@@
Abstract
The proliferation of generative models, such as Generative Adversarial Networks (GANs), Diffusion Models, and Variational Autoencoders (VAEs), has enabled the synthesis of high-quality multimedia data. However, these advancements have also raised significant concerns regarding adversarial attacks, unethical usage, and societal harm. Recognizing these challenges, researchers have increasingly focused on developing methodologies to detect synthesized data effectively, aiming to mitigate potential risks. Prior reviews have predominantly focused on deepfake detection and often overlook recent advancements in synthetic image forensics, particularly approaches that incorporate multimodal frameworks, reasoning-based detection, and training-free methodologies. To bridge this gap, this survey provides a comprehensive and up-to-date review of state-of-the-art techniques for detecting and classifying synthetic images generated by advanced generative AI models. The review systematically examines core detection paradigms, categorizes them into spatial-domain, frequency-domain, fingerprint-based, patch-based, training-free, and multimodal reasoning-based frameworks, and offers concise descriptions of their underlying principles. We further provide detailed comparative analyses of these methods on publicly available datasets to assess their generalizability, robustness, and interpretability. Finally, the survey highlights open challenges and future directions, emphasizing the potential of hybrid frameworks that combine the efficiency of training-free approaches with the semantic reasoning of multimodal models to advance trustworthy and explainable synthetic image forensics.
keywords:
@@ Abstract
highlights
Research highlights item 1
Research highlights item 2
Research highlights item 3
1 Introduction
@@
The advent of advanced generative models has enabled the creation of highly realistic synthetic images. These images are synthesized through various approaches, including conditional methods such as image-to-image translation and text-to-image translation, as well as unconditional generation. Generative models based on Generative Adversarial Networks (GANs) Goodfellow et al. (2014), Diffusion Models Sohl-Dickstein et al. (2015); Ho et al. (2020), Variational Autoencoders (VAEs) Kingma (2013), and Autoregressive Models Van der Maaten and Hinton (2008); Van Den Oord et al. (2016) dominate the current literature on image generation. Figure 1 provides a simple illustration of the architectures of these model families. While these four generative model families—GANs, Diffusion Models, VAEs, and Autoregressive—have dominated the landscape of image generation, other methods such as Normalizing Flows Rezende and Mohamed (2015) have also made significant contributions and deserve attention. With these advancements, several state-of-the-art generative models have been made publicly available alongside commercially accessible tools such as Adobe Firefly Adobe Inc. (2024), MidJourney MidJourney (2024), DALL·E 3 OpenAI (2024), and Imagen 3 Google DeepMind (2024).
@@
Although generative models have enabled advancements in image synthesis, their accessibility introduces critical concerns related to misinformation, privacy, and security. The ability to generate highly realistic AI-synthesized images raises ethical and security risks, including deceptive media Sandrini and Somogyi (2023), identity fraud Zhang et al. (2025), and geopolitical manipulation Mylrea (2025). Consequently, forensic detection has become an essential area of research, aiming to reliably distinguish AI-generated imagery from real-world photographs. Despite significant progress, existing detection techniques face challenges in generalizability, robustness, scalability, and reasoning across evolving generative architectures. The detection of images generated by both conditional and unconditional generative methods has emerged as a rapidly growing research area. Despite this increasing interest, a comprehensive review that consolidates and systematically categorizes existing detection methodologies is still lacking. While prior surveys have presented reviews on the detection of image manipulations achieved through traditional computer-based techniques Sencar and Memon (2013), image editing tools Zheng et al. (2019), and deepfake methodologies Wang et al. (2024), our focus diverges from these areas and addresses the limitations in this domain. This survey fills this gap by providing a structured analysis of detection approaches, classifying them based on their underlying techniques and contributions, and focusing on recent advances that incorporate multimodal techniques, reasoning with large language models, and training-free detectors for enhanced forensic accuracy. Given the increasing relevance of such methods in forensic image analysis, it is crucial to explore their long-term applicability and effectiveness in generative image detection.
Figure 1: Illustration of Simplified Architecture of GAN, Diffusion, VAE and Autoregressive Generative Models.
To establish a unified framework, this survey introduces a core taxonomy that organizes artificial intelligence-generated image (AIGI) detection methodologies into seven categories: spatial-domain analysis, frequency-domain analysis, fingerprint analysis, patch-based analysis, training-free methods, multimodal and reasoning-based models, and commercial detection frameworks. Each category represents a distinct perspective on identifying AI-generated imagery and serves as the foundation for the subsequent sections of this paper. By examining the evolving landscape of adversarial threats in generative AI, this study highlights key challenges, methodological trends, and emerging opportunities, reinforcing the need for continued research in forensic detection of AIGI.
2 Reviews
@@
In the present survey of detection methods for AI-generated images, we categorize them into seven distinct groups, with the final category comprising commercial methods. For each group, we discuss the core proposal of each method in ascending order based on publication date. At the end of each subsection, we present a table summarizing whether the experimental evaluations of the methods satisfy three key criteria, described as follows:
@@
3. Cross-Scene: This criterion assesses whether a method’s performance was tested across datasets with distinct data distributions. For instance, a detector trained on images from the ImageNet Russakovsky et al. (2015) dataset and evaluated on images from the LSUN-Bedroom Yu et al. (2015) dataset satisfies the cross-scene requirement. Importantly, methods that meet the cross-scene criterion typically also satisfy the cross-category requirement, although the converse is not necessarily true.
@@
Figure 2: Categorization of Detection Methods Based on Core Architecture and Methodological Proposals. The figure illustrates the division of detection methods into categories. Each category is highlighted using blue-colored rectangles. Some methods are connected to multiple categories, shown using light-red highlights and arrows pointing to the respective sub-categories.
2.1 Spatial Domain Analysis / Spatial Feature Analysis Methods
@@
2.1.1 CNN-Generated Images Are Surprisingly Easy to Spot: Wang et al. (2020)
Wang et al. Wang et al. (2020b) conducted a pivotal study demonstrating that a detection method trained on images generated by a single generative model, particularly a GAN, can generalize to detect synthetic images from a variety of unseen CNN-based generative models. This finding challenges the previously established view that cross-model generalization is inherently difficult for forensic classifiers. To evaluate this, the authors trained ProGAN Karras (2017) on the LSUN dataset Yu et al. (2015), producing a dataset of 720K images for training and 4K for validation, with an equal split between real and generated images. They employed a ResNet-50 He et al. (2016), pretrained on ImageNet, as a binary classifier. Robust feature learning was facilitated through various data augmentation strategies, including: (a) no augmentation, (b) Gaussian blur, (c) JPEG compression, and (d) combinations of both with varying probabilities (50% and 10%). The trained classifier demonstrated strong generalization capabilities, successfully detecting images synthesized by other prominent generative models such as StyleGAN Karras et al. (2019), BigGAN Brock (2018a), CycleGAN Zhu et al. (2017), StarGAN Choi et al. (2018), and GauGAN Park et al. (2019). This work highlights the existence of common artifacts across CNN-generated images, suggesting that classifiers can leverage these shared patterns for generalizable detection across different architectures and tasks.
2.1.2 Detection and Attribution of GAN Images: Goebel et al. (2020)
Goebel et al. Goebel et al. (2020) introduced a comprehensive framework to detect, attribute, and localize GAN-generated images through co-occurrence matrix analysis and deep learning. This method builds on insights from steganalysis, leveraging pixel-level co-occurrence features to identify artifacts introduced during image synthesis. The approach begins by computing co-occurrence matrices from the RGB channels of the input image in four orientations: horizontal, vertical, diagonal, and anti-diagonal. Each co-occurrence matrix captures a 256×256 histogram of pixel value pairs, normalized and stacked into a tensor of size $256\times 256\times 12$. The matrices are defined as:
@@
These features are then processed using a modified XceptionNet Chollet (2017), designed for three key tasks: 1. Binary Detection: Classifying images as real or GAN-generated. 2. Multi-Class Attribution: Identifying the GAN architecture (e.g., ProGAN, CycleGAN). 3. Localization: Generating heatmaps to identify manipulated regions through patch-based analysis.
Extensive experiments conducted on over 2.76 million images demonstrated the effectiveness of this method across various GAN models, including ProGAN, StyleGAN, CycleGAN, StarGAN, and SPADE/GauGAN. Additionally, t-SNE Van der Maaten and Hinton (2008) visualizations showed a clear separation between real and GAN-generated images, reinforcing the model’s interpretability and robustness against varying JPEG compression levels and patch sizes. This framework advances GAN forensics by integrating detection, attribution, and localization into a unified pipeline.
2.1.3 Estimating Artifact Similarity with Representation Learning: Li et al. (2022), GASE-Net
Li et al. Li et al. (2022) introduced GASE-Net, a framework designed to detect GAN-generated images by estimating artifact similarity. This method addresses challenges in cross-domain generalization and robustness against post-processing, using a two-stage approach: representation learning and representation comparison. In the representation learning stage, a ResNet-50 backbone is modified with instance normalization (IN) applied to the shallow layers to enhance feature extraction by filtering out instance-specific biases. This ensures that the learned representations remain invariant across different domains while retaining category-level distinctions. Feature maps from reference images are aggregated element-wise to form domain prototypes, which serve as robust representations of GAN or pristine image domains.
@@
2.1.4 Local Intrinsic Dimensionality Analysis: Lorenz et al. (2023), AdaptedMultiLID
Lorenz et al. Lorenz et al. (2023) proposed a framework utilizing the Multi-Local Intrinsic Dimensionality (multiLID) method for detecting diffusion-generated images. This approach builds upon the earlier work Lorenz et al. (2022) and demonstrates strong performance in distinguishing both diffusion and GAN-generated images. The method starts by extracting low-dimensional feature maps using an untrained ResNet18 model. MultiLID is then computed to capture local growth rates of feature densities within the latent space. The multiLID feature vector for each sample $x$ is defined as:
@@
2.1.5 AI-Generated Image Detection: Baraheem et al. (2023)
Baraheem et al. Baraheem and Nguyen (2023) proposed a framework to detect GAN-generated images using transfer learning on pretrained classifiers. The authors compiled a diverse dataset called Real and Synthetic Images (RSI), consisting of 48,000 images synthesized by 12 GAN architectures across tasks such as image-to-image, sketch-to-image, and text-to-image generation. EfficientNetB4 Tan and Le (2019) achieved the best detection performance after fine-tuning on the RSI dataset. The model’s architecture was modified by replacing the classifier head with layers for global average pooling, batch normalization, ReLU activation, dropout, and a sigmoid output. The training utilized the Adam optimizer with a batch size of 64, an initial learning rate of 0.001, and data augmentation techniques such as horizontal flips. To facilitate model explainability, the authors incorporated four Class Activation Map (CAM) methods, GradCAM Selvaraju et al. (2017), AblationCAM Ramaswamy et al. (2020), LayerCAM Jiang et al. (2021), and Faster ScoreCAM Wang et al. (2020a), to visualize the discriminative regions influencing classification decisions.
2.1.6 Diffusion Reconstruction Error (DIRE): Wang et al. (2023)
Wang et al. Wang et al. (2023a) proposed DIffusion Reconstruction Error (DIRE), a novel method to detect diffusion-generated images by leveraging reconstruction errors from pre-trained diffusion models. The method addresses the limitations of previous detectors, which struggled to generalize across different diffusion models. The key idea is that diffusion-generated images can be reconstructed more accurately by the pre-trained DDIM model Song et al. (2020) than real images. DIRE is defined as the $L_{1}$-norm of the difference between the input image $x_{0}$ and its reconstructed counterpart $x^{\prime}_{0}$:
@@
The process involves applying forward noise to the input image and then performing a reverse denoising process to generate the reconstruction. A ResNet-50 classifier is trained using binary cross-entropy loss on these DIRE representations to distinguish between real and generated images. The authors demonstrated that DIRE achieves state-of-the-art performance on their proposed DiffusionForensics dataset, which includes images generated by various diffusion models across multiple domains (e.g., LSUN-Bedroom Yu et al. (2015), ImageNet Russakovsky et al. (2015), and CelebA-HQ Karras (2017)). Extensive experiments showed that DIRE not only excels in detecting diffusion-generated images but also generalizes well to unseen models and maintains robustness under perturbations like Gaussian blur and JPEG compression.
2.1.7 Classification and Explainable Identification: Bird and Lotfi (2024)
Bird and Lotfi Bird and Lotfi (2024) introduced a framework for detecting AI-generated images using a large-scale dataset, CIFAKE, which is detailed in the datasets section. Their classification approach employs a Convolutional Neural Network (CNN) that processes images through stacked convolutional and pooling layers, followed by fully connected layers with a final sigmoid activation for binary classification. As seen in Baraheem and Nguyen (2023), the study emphasizes explainability by implementing Gradient Class Activation Mapping (Grad-CAM) Selvaraju et al. (2017), which generates heatmaps to highlight regions influencing the model’s decisions. Grad-CAM computes importance weights $\alpha_{k}$ for each feature map $A_{k}$, producing a visual explanation as: